# Exploring baby Julia sets in parameter space slices for Generalized McMullen Maps

### v1

@@ 7.2. Hyperbolic components
Remember in viewing the diagram of Figure 12 that $\hat{a}_{2j}$ is just one map, in the above picture $r_{n,\hat{a}_{2j}}(v_{-})$ stays put in its $D_{+}$, then as $a$ varies below, its image under $\tau$, which is $r_{n,a}(v_{-}(a))$, varies. Now, again thinking of the top half of the picture, $D_{+}(\hat{a}_{2j})$ is a quasi-disk (from Lemma 3.1), and for $r_{n,\hat{a}_{2j}},$ one critical point is fixed ($v_{+}$), and the other, $v_{-}$, must lie in a pre-image of $D_{+}$ because $\hat{a}_{2j}$ is near $a_{2j}$ in its hyperbolic component, and at $a_{2j}$, $v_{-}$ maps to $v_{+}$. So $D_{-}(\hat{a}_{2j})$ is also a quasi-disk.
Now as $a$ varies in $\mathcal{H}_{2j}^{*}$, $\tau_{a}^{2j}$ is a holomorphic function of $a$ (and injective in $z$), and $\phi_{a}$ depends continuously on $a$ for any $a$ in the interior of a hyperbolic component (by Lemma 3.1). We argue that $\Phi(a)$ sweeps out $\mathbb{D}\setminus\{0\}$ as $a$ sweeps out $\mathcal{H}_{2j}^{*}$. We have that $v_{-}$ is essentially $a^{1/2n}-4\sqrt{a}$. So what can cause a problem for $\Phi_{j}$? A problem with a square root or $1/2n$ root can happen for hyperbolic components which run into $a=0$, but we have excluded these with our restrictions on $j$. Or, another critical orbit interfering, but we are controlling this with our hypotheses on this subfamily: one critical orbit is fixed, the other is the only one which is independent.
Essentially, there are only two critical values, $v_{+}$ is stuck, we are in a hyperbolic component avoiding $a=0$, so $v_{-}$ can have its “full experience” as the parameter $a$ varies in $\mathbb{C}$ in a neighborhood of $a_{2j}$. We have a quasi-disk $D_{+}$ in dynamical space, and then it has a preimage under $r$ that’s a 2:1 branched covering from $D_{-}$ to $D_{+}$, so $D_{-}$ is a quasi-disk. And $r(v_{-})$ can be “anywhere” in $D_{+}$, as you vary $a$, so $v_{-}$ can be anywhere in $D_{-}$, so, $\phi_{a}(r(v_{-}))$ sweeps out the whole unit disk as $a$ varies in $\mathcal{H}_{2j}$. Now we define “anywhere” as relative to the corresponding points for the map $r_{n,\hat{a}_{2j}}$. That is, as $a$ varies in a neighborhood about $\hat{a}_{2j}$, a point sweeps out the $\mathbb{D}$ (in the upper right of the figure) which is the image under $\phi_{\hat{a}_{2j}}$ of $D_{+}(\hat{a}_{2j})$, thus the preimage under $\phi_{\hat{a}_{2j}}$ sweeps out the disk $D_{+}(\hat{a}_{2j})$, hence the preimage under $r_{n,\hat{a}_{2j}}$ sweeps out the disk $D_{-}(\hat{a}_{2j})$, and that is the reason you see a disk in parameter space. That gives $\Phi$ is surjective, and its injective since $\tau$ is a holomorphic motion and $\phi$ is a q.c.-homeomorphism, and it’s continuous because $\tau$ and $\phi$ vary continuously with $a$. This yields a copy of $\mathbb{D}$ in parameter space (since $v_{-}$ is just calculated from $a$ by simple complex arithmetic, and roots, and our hyperbolic components avoid 0).
@@ 7.3. Future work on the location of the topological disk components
To establish more specific results about these topological disk components for $r_{n,a}$, such as defining explicit regions for each $n$ in which they are contained, one could first attempt to prove Conjecture 2 in general then apply to this family, or vice-versa establish this in the simpler case first to point the way to the future work of generalization.
@@
For this family of maps $r_{n,a}$, one could try to define some regions $\mathscr{A}_{n}$ in parameter space, then in terms of $n$ calculate or bound their distance from 0 (where $D_{-}=D_{+}$)). The regions could be defined dynamically or by simpler polar rectangles. Then the approach would be to show that for the boundary of these regions, or at least for one of these regions for a start, $r(v_{-})$ loops around the boundary of $D_{+}$.
Looking at the pictures (see Figure 4), if $n$ is even there’s a topological disk centered on the positive real axis (in dynamical plane), in a region $\mathscr{A}_{n}$ if $n$ is odd there are two symmetrical ones in quadrants 1 and 4, the indices are $n\pm 1$}.
The “worst” case is $n=3$ in which $\mathscr{A}_{2}$ and $\mathscr{A}_{4}$ are the only disks. A starting approach could be to start with larger $n$ to make the argument easier, then consider decreasing $n$.
Perhaps one can show that $\mathscr{A}_{n}$ or $\mathscr{A}_{n\pm 1}$ is outside of a lower radius and then say in the right half plane. (For $n=3$ this is not quite true, the cardioid has cusp at 0 and the two disks look centered along the imaginary axis).
We don’t know how big the topological disks or baby Mandelbrot’s might be, but we have the centers.
@@
One could try using Lemma 4.2 or Corollary 4.3, for any $\varepsilon>0$, there is an $n$ sufficiently large that, if $l\leq 1/32$ and $u>3/8$, then $M_{n}(r_{n,a})\subset\mathbb{A}(l,u)+1/8\subset\mathbb{D}(0,1/2+\varepsilon).$ In proof of the Lemma 4.2, we also showed if $a\in\mathbb{D}(1/8,1/32)$ (i.e., |a1/8|<1/32|a-1/8|<1/32) then $|v_{-}|\leq L(\infty)<0.614,$ and if $|a|>1/2$ then for sufficiently large $n$, $|v_{-}|\geq 1+\varepsilon$, for $\varepsilon$ depending on how for $u$ is from $3/8$ (maybe not same $\varepsilon$ as we stated above about the disk though we doubt it matters). One could start with an annulus and chop off say left of the line $\Re(z)=1/8$, try to establish topological disk components to the right (one or two depending on if $n$ is even or odd).
One must avoid $a$ close to 0, for which $v_{-}$ too close to $v_{+}$. One needs the Fatou components of $v_{-}$ and $v_{+}$ staying disjoint, to avoid the degenerate case discussed near the beginning of Section 3.
@@
Remember $\operatorname{Arg}$ has range $(-\pi,\pi]$. And $a_{k}$ is the parameter such that $v_{-}=w_{k}$ the $k^{th}$ critical point. Note $w_{2j}$ is not the critical point of $r$ for any $a\in W_{2j}$ it’s the specific critical point of $r$ for $a_{2j}$ specifically.
@@
So a strategy could be: for any $a\in\mathscr{A}_{2j}$ (with center $a_{2j}$ at which $v_{-}=w_{2j}$), for $j\in\{1,\ldots n-1\},$ (so excluding $a_{0}=0$), find bounds in terms of $n$ for the location of $a$ and/or $v_{-}$ and/or $w_{2j}$. From definition of $W_{2j}$, we know for every $a$ in the interior of $W_{2j}$, we have $\frac{|a|^{1/n}}{2}<|v_{-}|<2$ and $|\operatorname{Arg}(v_{-})-\operatorname{Arg}(w_{2j})|<\frac{\pi}{2n}.$
Depending on the definition of the $W_{k}$’s one may need to prove the $W$’s partition up a region containing the boundedness locus in a useful way.

### v2

@@ 7.2. Hyperbolic components
Remember in viewing the diagram of Figure 12 that $r_{n,\hat{a}_{2j}}$ is just one map, and in the above picture $r_{n,\hat{a}_{2j}}(v_{-})$ stays put in its $D_{+}$, then as $a$ varies below, its image under $\tau$, which is $r_{n,a}(v_{-}(a))$, varies. Again thinking of the top half of the picture, $D_{+}(\hat{a}_{2j})$ is a quasi-disk as shown in Lemma 3.1, and for $r_{n,\hat{a}_{2j}},$ one critical point, $v_{+}$, is fixed, and the other, $v_{-}$, must lie in a pre-image of $D_{+}$, because $\hat{a}_{2j}$ is near $a_{2j}$ in its hyperbolic component, and at $a_{2j}$, $v_{-}$ maps to $v_{+}$. So $D_{-}(\hat{a}_{2j})$ is also a quasi-disk.
Now, as $a$ varies in $\mathcal{H}_{2j}^{*}$, $\tau_{a}^{2j}$ is a holomorphic function of $a$ (and injective in $z$), and $\phi_{a}$ depends continuously on $a$ for any $a$ in the interior of a hyperbolic component by Lemma 3.1. We argue that $\Phi(a)$ sweeps out $\mathbb{D}\setminus\{0\}$ as $a$ sweeps out $\mathcal{H}_{2j}^{*}$. We have that $v_{-}$ is essentially $a^{1/2n}-4\sqrt{a}$. One theoretical problem for $\Phi_{j}$ could be a problem with a square root or $1/2n$ root, which can happen for hyperbolic components which run into $a=0$, but we have excluded these with our restrictions on $j$. Or, another problem would be a critical orbit interfering, but we are controlling this with our hypotheses on this subfamily that one critical orbit is fixed and the other is the only one which is independent.
Essentially, since there are only two critical values, $v_{+}$ is stuck in a hyperbolic component avoiding $a=0$, so $v_{-}$ can have its “full experience” as the parameter $a$ varies in $\mathbb{C}$ in a neighborhood of $a_{2j}$. We have a quasi-disk $D_{+}$ in dynamical space, and which has a preimage $D_{-}$ on which $r$ is a 2:1 branched covering to $D_{+}$, so $D_{-}$ is a quasi-disk. Since $r(v_{-})$ can be “anywhere” in $D_{+}$ as one varies $a$, we see that $v_{-}$ can be anywhere in $D_{-}$, so, $\phi_{a}(r(v_{-}))$ sweeps out the whole unit disk as $a$ varies in $\mathcal{H}_{2j}$. Now we define “anywhere” as relative to the corresponding points for the map $r_{n,\hat{a}_{2j}}$. That is, as $a$ varies in a neighborhood about $\hat{a}_{2j}$, a point sweeps out the $\mathbb{D}$ (in the upper right of the figure) which is the image under $\phi_{\hat{a}_{2j}}$ of $D_{+}(\hat{a}_{2j})$, thus the preimage under $\phi_{\hat{a}_{2j}}$ sweeps out the disk $D_{+}(\hat{a}_{2j})$, hence the preimage under $r_{n,\hat{a}_{2j}}$ sweeps out the disk $D_{-}(\hat{a}_{2j})$, and that is the reason you see a disk in parameter space. That gives $\Phi$ is surjective, and its injective since $\tau$ is a holomorphic motion and $\phi$ is a q.c.-homeomorphism, and it’s continuous because $\tau$ and $\phi$ vary continuously with $a$. This yields a copy of $\mathbb{D}$ in parameter space, since $v_{-}$ is just calculated from $a$ by simple complex arithmetic and roots, and our hyperbolic components avoid 0.
@@ 7.3. Future work on the location of the topological disk components
To establish more specific results about these topological disk components for $r_{n,a}$, such as defining explicit regions for each $n$ in which they are contained, one could first attempt to prove Conjecture 2 in general then apply it to this family, or vice versa, to establish this conjecture in the simpler case first to point the way to the future work of generalization.
@@
For this family of maps $r_{n,a}$, one could try to define some regions $\mathscr{A}_{n}$ in parameter space, for which one could calculate or bound their distance from 0 in terms of $n$, since 0 is to be avoided as it is where $D_{-}=D_{+}$. The regions could be defined dynamically or more simply using polar rectangles. The approach would be to show that for the boundary of these regions, or at least for one of these regions as a start, $r(v_{-})$ loops around the boundary of $D_{+}$.
Considering computer generated images such as those provided in Figure 4, one can observe that if $n$ is even there’s a topological disk centered on the positive real axis in dynamical plane, in a region we denote $\mathscr{A}_{n}$, and if $n$ is odd, there are two symmetrical ones in quadrants 1 and 4, which would have the indices $n\pm 1$}. Given this line of thinking, the “worst” case is $n=3$ in which $\mathscr{A}_{2}$ and $\mathscr{A}_{4}$ are the only disks. A starting approach could be to start with larger $n$ to make the argument easier, then consider decreasing $n$. Perhaps one can show that $\mathscr{A}_{n}$ or $\mathscr{A}_{n\pm 1}$ is outside of a lower bound on radius and then say in the right half plane, although this would not be true for $n=3$, where the cardioid has cusp at 0 and the two disks look centered along the imaginary axis. As of now, we don’t know how big the topological disks or baby Mandelbrot’s might be, but we do have the centers.
@@
One could try using Lemma 4.2 or Corollary 4.3, showing that for any $\varepsilon>0$, there is an $n$ sufficiently large that, if $l\leq 1/32$ and $u>3/8$, then $M_{n}(r_{n,a})\subset\mathbb{A}(l,u)+1/8\subset\mathbb{D}(0,1/2+\varepsilon).$ In proof of the Lemma 4.2, we also showed if $a\in\mathbb{D}(1/8,1/32)$ (i.e., |a1/8|<1/32|a-1/8|<1/32) then $|v_{-}|\leq L(\infty)<0.614,$ and if $|a|>1/2$, then for sufficiently large $n$, $|v_{-}|\geq 1+\varepsilon_{1}$, for $\varepsilon_{1}$ depending on how far $u$ is from $3/8$. One could start with an annulus and try to establish one or two (depending on the parity of $n$) topological disk components to the right of the line $\Re(z)=1/8$. The reason for staying to the right of the line is that one must avoid $a$ close to 0, for which $v_{-}$ too close to $v_{+}$. One needs the Fatou components of $v_{-}$ and $v_{+}$ staying disjoint, to avoid the degenerate case discussed near the beginning of Section 3.
@@
Furthermore, remember $\operatorname{Arg}$ has range $(-\pi,\pi]$, and $a_{k}$ is the parameter such that $v_{-}=w_{k}$ the $k^{th}$ critical point. Note $w_{2j}$ is not the critical point of $r$ for any $a\in W_{2j}$, but the specific critical point of $r$ for the hyperbolic center $a_{2j}$.
@@
A strategy could be to find bounds in terms of $n$ for the location of $a$, $v_{-}$ and/or $w_{2j}$, for any $a\in\mathscr{A}_{2j}$ (with center $a_{2j}$ at which $v_{-}=w_{2j}$), for $j\in\{1,\ldots n-1\},$ (so excluding $a_{0}=0$). As a start, from definition of $W_{2j}$, we know for every $a$ in the interior of $W_{2j}$, we have $\frac{|a|^{1/n}}{2}<|v_{-}|<2$ and $|\operatorname{Arg}(v_{-})-\operatorname{Arg}(w_{2j})|<\frac{\pi}{2n}.$ Depending on the definition of the $W_{k}$’s one may need to prove that the $W$’s partition up a region containing the boundedness locus in a useful way.